ABC Ltd., a technology company, issues a $47.2 million IPO providing proceeds to ABC of $5.31 per share, from an offer price to the public of $5.9 per share. The company's legal fees, ASIC registration fees, and other administrative costs are $164,000. The company's share price increases 5.6 per cent on the first day. What is the total amount of the underwriter’s spread (in millions of dollars to the nearest three decimal places; don’t use the $ sign eg 7.897)?
Answer #1

Gross Issue = $ 47.2 million, Net Proceeds = $ 5.31 per share and Offer Price = $ 5.9

Number of Shares Sold = 47.2 / 5.9 = 8 million

Net Proceeds = 5.31 x 8 = $ 42.48 million

IPO Costs (ASIC fees + Legal Fees + Administrative Costs) = $ 164000

Therefore, Total Amount of Underwriter's Spread = (47.2 - 42.48 - 0.164) = $ 4.556 million